OpenStudy (anonymous):

Solve 2x^2 + 8x +20= 0 by completing the square.

OpenStudy (anonymous):

I've started it but I need help...

OpenStudy (anonymous):

so I know you have to divide the 8x by 2, which is 4 and then you square 4 to get 16. then you add 16 to both sides so now your equation is 2x^2+8x+16=-4, but this is where I'm stuck - with the factoring

OpenStudy (anonymous):

make the coefficient of x^2 as 1

OpenStudy (anonymous):

how would that help?

OpenStudy (anonymous):

can someone please help me!

OpenStudy (anonymous):

i just need help factoring

OpenStudy (anonymous):

x^2+4x+10=0 or x^2+4x=-10 add both sides (4/2)^2 i.e.,4 x^2+4x+4=-10+4 (x+2)^2=-6=6*-1=6*i^2 (x+2)^2-6i^2=0 (x+2)^2-(i sqrt 6)^2=0 (x+2+i sqrt 6)(x+2- i sqrt 6 )=0